Let u=(-1/2,-1,2/3) ,v=(4,-1,-3) and θ=1/2. Find

i) u.θv

ii)A non-zero vector orthogonal to both u and v


1
Expert's answer
2021-08-12T14:59:31-0400

u=(-1/2,-1,2/3)

v=(4,-1,-3)


u.v = ?

θ = ?


(i)

"u = -1/2i -j + 2/3k\\
|u| =\sqrt{(-1/2)²+( -1)²+ (2/3)²) }\\=1.302\\"


"v = 4i-j -k\\
|v|= \sqrt{(4)²+( -1)²+ (-3)²)} =5.10"


dot product (u.v) = "(a_ib_i + a_jb_j + a_kb_k)"

"= ((-1/2×4)+ (-1×-1) +(2/3×-3))\\
= (-2+1-2) = -3"


"\theta=\cos^{-1} \dfrac{u.v}{|u||v|} =\cos^{-1}\dfrac{-3}{1.302× 5.10}\\"


"=\cos^{-1} -0.25 = 116.9°"


(ii)

"u = -1/2i -j + 2/3k\\
v = 4i-j -3k"

A vector "C \neq O" that is orthogonal to both.


Since "C = (c_1, c_2, c_3)" must be orthogonal to both u and v;

"\begin{aligned} u\cdot C &= 0 & \implies & &-1/2c_1 - c_2 + 2/3c_3 &= 0 \\ v\cdot C &= 0 & \implies & &4c_1 -c_2 - 3c_3 = 0 \end{aligned}"


From the first equation; "c_2= -1\/2c_1 +2\/3c_3"


"[ 4c_1 -c_2 - 3c_3 = 0 \quad \implies \quad 4c_1 -(-1/2c_1 + 2/3c_3) - 3c_3= 0 \quad \implies \\
4c_1+1/2c_1-2/3c_3-3c_3=0\\ \implies 9/2\ c_1 = 11/3\ c_3"

"\therefore c_1 = \dfrac{22}{27}\ c_3"


Since "c_3" is then arbitrary we choose "c_3 = 27" and obtain

"[ c_2 = 7 \qquad c_1 = 22]"


Therefore, "C = (22,7,27)"